6 minutes. For each age group shown in the table, the median distance walked Two hundred and eighty boys completed a test that measures the distance that the boy can walk on a flat, hard surface by the boys that age group is given. Median Six-Minute Walk Distance (meters) Representative Age (Midpoint of Age Group) Age Group 3-5 4.0 543.3 585.0 6-8 7.0 9-11 10.0 666.3 12-15 13.5 702.1 17.0 16-18 728.6
(c) Calculate the five residuals, make sure to use at least 5 decimals for the slope and intercept when calculating the predicted values and residuals. (Round your answers to three decimal places.) Representative Age (x) Residual 4.0 -8.54 7,0 -12.141 26.869 10.0 10.04 13.5 -16.174 17.0
Answer #1
X Y (x-x̅)² (y-ȳ)² (x-x̅)(y-ȳ)
4 543.3 39.69 10355.10 641.09
7 585 10.89 3607.20 198.20
10 666.3 0.09 451.14 -6.37
13.5 702.1 10.24 3253.56 182.53
17 728.6 44.89 6978.93 559.72
ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 51.5 3225.3 105.8 24645.9 1575.16
mean 10.30 645.06 SSxx SSyy SSxy

sample size ,   n =   5          
here, x̅ = Σx / n=   10.30   ,     ȳ = Σy/n =   645.06  
                  
SSxx =    Σ(x-x̅)² =    105.8000          
SSxy=   Σ(x-x̅)(y-ȳ) =   1575.2          
                  
estimated slope , ß1 = SSxy/SSxx =   1575.2   /   105.800   =   14.8881
                  
intercept,   ß0 = y̅-ß1* x̄ =   491.7127          
                  
so, regression line is   Ŷ =   491.712665   +   14.888091   *x

c)

X residual,ei=y-yhat
4 551.27 -7.965
7 595.93 -10.929
10 640.59 25.706
13.5 692.70 9.398
17 744.81 -16.210